OpenStudy (love_to_love_you):

PLEASE HELP WILL FAN AND MEDAL AND TESTIMONY Write each polynomial function in standard form, classify it by degree, and determine the end behavior of its graph. Show all work. 1. y = 12-x^4 2. y = 2x^3 - 6x + 3x^2 - x^4 + 12 3. y = 10 - 3x^3 + 3x^2 + x^4

OpenStudy (love_to_love_you):

@mathmath333

OpenStudy (mathmath333):

\(\large \begin{align} \color{black}{ 1.)~~ y = 12-x^4 \hspace{.33em}\\~\\ \normalsize \text{notice the power of x } \hspace{.33em}\\~\\ y = 12-x^{\color{red}{4}} \hspace{.33em}\\~\\ \normalsize \text{so its 4 degree monomial as variable x is only once , } \hspace{.33em}\\~\\ \normalsize \text{similarly try out others } \hspace{.33em}\\~\\ }\end{align}\)

OpenStudy (mathmath333):

\(\large \begin{align} \color{black}{ 1.)~~ y = 12-x^4 \hspace{.33em}\\~\\ \normalsize \text{this can be written as } \hspace{.33em}\\~\\ y = 12-(x^{2})^2 \hspace{.33em}\\~\\ \normalsize \text{so the parent function here is } \hspace{.33em}\\~\\ y=-x^2 \hspace{.33em}\\~\\ \normalsize \text{so } \hspace{.33em}\\~\\ y-12 = -(x^{2})^2 \hspace{.33em}\\~\\ \normalsize \text{y intecept is 12, so } \hspace{.33em}\\~\\ }\end{align}\)

OpenStudy (mathmath333):



OpenStudy (mathmath333):